### v1

# Binaural Target Speaker Extraction using HRTFs and a Complex-Valued Neural Network

Abstract
In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers. To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker. Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets in different languages. We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals. This deviates from conventional approaches that use spectrograms or treat the real and imaginary components of the STFT as separate real-valued inputs. We begin by evaluating the method in an anechoic, noise-free scenario, where it demonstrates excellent extraction performance while effectively preserving the binaural cues of the target signal. Next, we test a modified variant under mild reverberation conditions. This version remains robust in reverberant conditions, maintaining speech clarity, preserving source directionality, and simultaneously reducing reverberation. Demo-page: https://bi-ctse-hrtf.github.io
Index Terms:
Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
I Introduction
Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking. Such a scenario is often referred to as the cocktail party problem [1]. While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments. Binaural Target Speaker Extraction (Bi-TSE) aims at selectively extracting one desired speaker of the multiple competing speakers, in a binaural setting. In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3]. Most hearing aids feature multiple microphones on each earpiece and fall into one of the two categories: 1) Bilateral systems, in which each device operates independently, yielding a single-channel output. 2) Shared systems, where both earpieces share all microphones, producing a dual-channel output.
Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears. In bilateral systems, where each device processes its own channel independently, it is particularly challenging to preserve these binaural cues. In contrast, Multi input Multi output (MIMO) solutions jointly process both channels, yielding more natural and spatially accurate outputs. For example, [4] presents a Binaural Linearly Constrained Minimum Variance (BLCMV) beamformer that separates sources while maintaining binaural cues. More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
Unlike Blind Source Separation (BSS) algorithms, which separate all sources, Target Speaker Extraction (TSE) algorithms leverage prior knowledge of the target speaker(s) to isolate their voice. The prior knowledge about the speaker—commonly referred to as “clue” [6]—can include any additional information about the target speaker. For instance, it might be visual data depicting the speaker’s appearance as in [7] that uses a face image to extract its corrospondig speech, or [8] that uses both audio and visual clues, where the audio clue is an enrollment utterance and the visual clue is a video of the speaker speaking within the mixture. Another popular clue is a pre-recorded sample used to extract speaker-specific features, as in [9, 10, 11, 12].
While examining prior works, such as [13], we observed the particular challenge of separating two voices with highly similar characteristics—for example, speakers of the same gender. Consequently, we opted for an approach that does not depend on speaker–embedding–based clues. Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better. For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA. These spatially guided methods consistently outperform single-microphone solutions.
In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that leverages the listener’s HRTF, aligned with the target speaker’s DOA, as an extraction clue. Our approach employs a complex‑valued DNN that operates directly in the STFT domain. Unlike real‑valued models that treat real and imaginary parts separately—or architectures that rely solely on the magnitude spectrogram—our network processes the full complex representation, more effectively preserving phase information. The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved. Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics. The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases.
II Problem Formulation
We address the problem of multiple concurrent speakers, captured by two microphones, each mounted on one of the listener’s ears in a noiseless environment. For clarity, we focus on the simpler case of two concurrent speakers, as formulated below.
Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation). Similarly, the interfering speaker’s DOA is ($\theta_{i}$,$\phi_{i}$). Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively. This signal can be written as
$\bm{x}_{b}(n,k)\;=\;s_{d}(n,k)\,\bm{h}(\theta_{d},\phi_{d},k)+s_{i}(n,k)\,\bm{h}(\theta_{i},\phi_{i},k)$ (1)
where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech of the desired speaker and the interfering speaker, respectively. Let $\gamma=\{d,i\}$ denote the speaker index, either the desired ($d$) or the interferer ($i$). Here, $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ is a Transfer Function (TrF) modeling acoustic propagation from source $\gamma$ to both earpieces as a superposition of HRTFs associated with the DOAs of all reflections. The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, represents the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$. According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA. Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity. Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$ Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
Our work has three main objectives: 1) to extract the target speaker at a known DOA from binaural input using the listener’s HRTF; 2) to generate a binaural output that preserves the spatial cues of the desired source; and 3) to dereverberate the output, retaining only the direct-path HRTF.
III Proposed Method
We propose a deep neural network model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task. Given the HRTF from the desired source DOA $(\theta_{d},\phi_{d})$, namely $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$. We denote the extracted signal as $\hat{\tilde{\bm{s}}}_{d}$.
III-A Architecture
The proposed model architecture is based on the approach presented in [14], modified to fully handle complex-valued inputs, rather than the real and imaginary components independently.
Encoder-Decoder Blocks of the U-net: The encoder comprises five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation function:
$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\left(\mathcal{R}(z)\right)+\mathrm{j}\cdot\text{ReLU}\left(\mathcal{I}(z)\right).$ (2)
Here $\text{ReLU}(x)=\max(0,x)$ and $\mathcal{R}(z),\mathcal{I}(z)$ denote the real and imaginary parts of z, respectively. Additionally, $\mathrm{j}$ represents the imaginary unit. Similarly, the decoder comprises five up-convolution blocks, each employing transposed convolutional layers. Skip connections are utilized between corresponding encoder and decoder blocks of the U-net structure to improve reconstruction accuracy. Additionally, the decoder includes a final convolutional layer with no activation function, enabling the representation of both positive and negative values at the output. This design choice is essential since the STFT representation inherently includes both positive and negative real and imaginary components.
Bottelneck Processing: The extraction clue is applied to the encoded input features at the bottleneck, with self-attention layers employed both before and after its application. However, unlike traditional methods, we implement these self-attention layers directly on complex-valued tensors by replacing the softmax activation function with the complex-valued $\mathbb{C}\text{ReLU}$.
To capture cross-channel (left and right) information in the HRTF, we apply the self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage. A block diagram of this architecture is presented in Fig. 1.
[…]
Figure 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed signal at the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of estimated desired signal.
III-B Loss function
We initially trained the proposed model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
[…]
where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all samples of the target and estimated signals, respectively. However, relying solely on the SI-SDR loss proved challenging in preserving accurate spectral details, resulting in noticeable musical noise artifacts in the extracted signals. Musical noise refers to tonal distortions that occur due to inaccuracies or rapid variations in the estimated signal’s frequency-domain representation [17]. Such artifacts are commonly observed in signal enhancement tasks when spectral continuity and consistency are not adequately enforced. To mitigate these artifacts, we introduced an additional frequency-domain loss, the $L1$- Mean Absolute Error (MAE) loss:
$\mathcal{L}_{\text{MAE}}=\frac{1}{2KN}\sum_{c=\{l,r\}}\sum_{k=0}^{K-1}\sum_{n=0}^{N-1}\left||\tilde{s}_{d}^{c}(n,k)|-|\hat{\tilde{s}}_{d}^{c}(n,k)|\right|.$ (4)
Here, $c=\{l,r\}$ denotes the channels, left or right. To improve data utilization, both speakers were extracted from each mixed signal in each training step by alternating the HRTF clues and averaging the SI-SDR losses. The overall training loss is therefore defined in (5), as a weighted combination of the SI-SDR and $L1$-MAE losses:
$\mathcal{L}=\frac{1}{2}\sum_{\gamma=\{d,i\}}\left(-\text{SI-SDR}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})+\alpha\cdot\mathcal{L}_{\text{MAE}}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})\right).$ (5)
Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a two-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$). Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality.
IV Experimental Study
In this section, we describe the data-generation process and experimental setup, evaluate the performance of the proposed approach, and compare it with previous TSE methods.
IV-A Data Generation
Binaural Impulse Rsponses: The binaural Impulse Responses are formed by superimposing attenuated and time-shifted HRIRs, corresponding to acoustic reflections generated via the image method [18], as expressed in (6).
$\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)=\sum_{\ell=0}^{L}a_{\ell}\cdot\bar{\bm{h}}_{\text{hrir}}(\theta_{\ell},\phi_{\ell},t-\tau_{\ell})$ (6)
where $L$ is the reflection order. The direct path corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$.
The RIEC HRTF Dataset [19, 20]: A measured HRTF database consists of 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
-90
-80
-70
-60
-50
-40
-30
-20
-10
0
10
20
30
40
50
60
70
80
90
Refer to captionRefer to captionRefer to caption
Figure 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$ [^2] Images sourced from: https://www.freepik.com
We employed the SofaMyRoom framework [23] to simulate reverberant and anechoic conditions, similar to (6). This tool places a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a target reverberation time. It also supports zero-order simulations for anechoic scenarios. For training, we sampled $T_{60}\sim\mathcal{U}[0.2,0.8]\,$s, and for testing, $T_{60}\sim\mathcal{U}[0.1,0.4]\,$s. Speaker and listener positions were also drawn uniformly within the room’s dimensions. The simulation arrangement is illustrated in Fig. 2.
Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [24]. To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
IV-B Algorithm Setup
The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, and transformed using a 512-sample STFT with 75% overlap. Our model was trained in a supervised manner. For the anechoic setting, the target signals were generated by convolving the desired speech signal with the corresponding HRTF. For the reverberant setting, the targets were obtained by simulating a reverberant room and using only the first arrival (i.e., zero-order reflection). The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and transformed to the frequency domain using Fast Fourier Transform (FFT).
IV-C Metrics
To evaluate our method’s performance, we used SI-SDR[27], Perceptual Evaluation of Speech Quality (PESQ)[28], and Short-time Objective Intelligibility (STOI)[29] as measures of speech quality and intelligibility. lFurthermore, to verify that the estimated signals retain their binaural characteristics, we compare the Probability Density Function (p.d.f.) of the ILD and ITD of the target signals and the extracted signals. The similarity of the p.d.f.s indicates that the spatial cues are preserved.
IV-D Competing Methods
To ensure a fair comparison, all baseline methods operate under an identical setup, namely, binaural TSE using only two microphones (one per ear). Accordingly, we compare our method against the BLCMV beamformer [4] and a recent DNN-based model [30], both designed specifically for this setup. In addition, we trained two variants of our model. The first is Bi-TSE-HRTF-RI, a real-valued model, in which the complex-value input features are split into Real-Imaginary (RI) components and concatenated, resulting in four input channels instead of the two input channels in the original scheme. The same procedure is applied to the HRTF clue. The second is Bi-TSE-DOA, a complex-valued model that incorporates the DOA as a clue instead of the HRTF. Here, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs and passed through a fully-connected layer to match the dimensionality of the bottleneck. For the specified variants, both terms of (5) remain intact, i.e., the target signal is the binaural rendering of the desired speech signal with the correct HRTF.
(a) SI-SDR[dB]
(b) PESQ
(c) STOI
Figure 3: Histograms of speech quality and intelligibility measures, before and after the application of the proposed TSE algorithm for the anechoic setting.
TABLE I: The evaluation datasets include 3k speech samples mixed with SIR sampled from $\mathcal{U}[0,5]$ dB. Speaker overlap is either sampled from $\mathcal{U}[0\%,100\%]$ (designated by) or set to 100% (designated by∗∗), with audio lengths of 5 s and 8 s, respectively. Best scores per metric and dataset are highlighted in bold.
Method Dataset Reverberant SI-SDR (dB) $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ $\Delta$ILD [dB] $\downarrow$ $\Delta$ITD [mS] $\downarrow$
Unprocessed LS∗∗ 2.4 1.36 0.78 - -
BLCMV Beamformer [4] LS∗∗ 0.86 3.93 0.95 0.7341 0.6032
Bi-CSim-TSE [30] LS 18.52 3.05 0.96 - -
Bi-TSE-HRTF (proposed) LS∗∗ 19.35 3.27 0.98 0.1311 0.0033
Unprocessed MLS-French∗∗ 2.37 1.29 0.77 - -
BLCMV Beamformer [4] MLS-French∗∗ 0.98 3.92 0.95 0.7092 0.5829
Bi-TSE-HRTF (proposed) MLS-French∗∗ 18.9 3.23 0.98 0.2697 0.0120
Unprocessed WSJ0∗∗ 0.06 1.23 0.73 - -
BLCMV Beamformer [4] WSJ0∗∗ 1.51 3.98 0.96 0.733 0.5669
Bi-TSE-DOA WSJ0∗∗ 16.3 2.47 0.97 0.1304 1.0145
Bi-TSE-HRTF-RI WSJ0∗∗ 21.36 2.97 0.99 0.0934 0.00435
Bi-TSE-HRTF (proposed) WSJ0∗∗ 20.37 3.49 0.99 0.1003 0.0027
Unprocessed WSJ0∗∗ -1.15 1.21 0.7 - -
BLCMV Beamformer [4] WSJ0∗∗ 1.37 1.6 0.82 0.3303 0.5439
Bi-TSE-DOA WSJ0∗∗ 9.53 2.08 0.93 0.2410 1.0125
Bi-TSE-HRTF (proposed) WSJ0∗∗ 10.37 2.60 0.95 0.1998 0.0036
IV-E Results
In Fig. 3, we show histograms of the SI-SDR, PESQ, and STOI metrics before and after applying our TSE algorithm in the anechoic scenario. All three distributions shift markedly to the right, demonstrating substantial gains across signal quality, perceptual quality, and intelligibility. This confirms the algorithm’s effectiveness in isolating the desired speaker.
In Table I, we compare our method against competing algorithms across all datasets under both anechoic and reverberant conditions. Our approach achieves significant gains on every metric except PESQ, where the BLCMV beamformer slightly outperforms the proposed method in anechoic case. This may be attributed to its distortionless response design criterion.
When comparing the two additional variants of our model, denoted as Bi-TSE-DOA and Bi-TSE-HRTF-RI in Table I, we observe that, in contrast to the HRTF-guided model, the DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and is unable to preserve the binaural cues. We evaluated this model under reverberant conditions to emphasize that the HRTF is intrinsically tied to our problem formulation. The superior performance of the HRTF-guided network supports these findings. On the other end, the RI model achieves comparable SI-SDR and binaural cue preservation to the complex-valued model but suffers from lower output quality, as shown by its lower PESQ score in the anechoic setting.
Remarkably, our approach demonstrates cross-dataset generalizability: although our model was trained exclusively on WSJ0-CSR1, it performs exceptionally well on other corpora, including a French dataset. We attribute this generalization to the absence of speaker-specific priors in our approach.
Figure 4: A polar plot showing the SI-SDR of the output signals for each DOA, with the desired speaker located at $\theta_{d}=-75^{\circ}$ (red) and the interfering speaker at $\theta_{i}=-10^{\circ}$ (blue). The bold circle represents the 0 dB SI-SDR threshold.
When inspecting the extracted signals, we found that the model effectively enhances the target speaker while nearly eliminating the interferer. To quantify its spatial resolution, we devised a “scanning” procedure: we iteratively extract signals using HRTFs corresponding to azimuth angles in $[-90^{\circ},90^{\circ}]$ at a fixed elevation angle and compute the SI-SDR of each extraction against the true target. Plotting these values yields a beam-pattern-like diagram (Fig. 4), which shows a pronounced peak at the correct DOA. Additionally, the width of the main lobe indicates that the model remains effective even when the HRTF clue deviates slightly from the true HRTF. The proposed method’s ability to preserve binaural cues is also evaluated and compared with the BLCMV, where applicable. Table I presents the mean ILD and ITD deviations. Figure 5 depicts the p.d.f.s of both cues for the target and extracted signals, showing a close match and confirming faithful spatial cue preservation.
(a) Mixture
(b) $\hat{\tilde{\bm{s}}}_{d}$
(c) $\hat{\tilde{\bm{s}}}_{i}$
Figure 5: The p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=90^{\circ}$, $\theta_{i}=15^{\circ}$, and $\phi_{d,i}=0^{\circ}$. Graphs produced by [31].
V Conclusions
This paper presents a speaker‑independent TSE method that leverages the listener’s HRTFs in both anechoic and reverberant settings, resulting in a personalized solution for the proposed problem. Through cross‑dataset evaluations, we demonstrate its robust generalization across diverse corpora. Our approach not only delivers high‑quality speech but also faithfully preserves the binaural character of the signal. Comparisons with other binaural TSE techniques underscore the importance of spatial cues tied to the listener’s HRTF for accurate and consistent target‑speaker extraction under varying acoustic conditions.

### v2

@@
-Binaural Target Speaker Extraction using HRTFs and a Complex-Valued Neural Network
+Binaural Target Speaker Extraction using HRTFs
 Abstract
@@
 To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker.
-Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets in different languages.
-We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals.
-This deviates from conventional approaches that use spectrograms or treat the real and imaginary components of the STFT as separate real-valued inputs.
-We begin by evaluating the method in an anechoic, noise-free scenario, where it demonstrates excellent extraction performance while effectively preserving the binaural cues of the target signal.
-Next, we test a modified variant under mild reverberation conditions.
-This version remains robust in reverberant conditions, maintaining speech clarity, preserving source directionality, and simultaneously reducing reverberation.
+Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets and languages.
+We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals, and compare it to a Real-Imaginary (RI)-based neural network, demonstrating the advantages of the former.
+We first evaluate the method in an anechoic, noise-free scenario, achieving excellent extraction performance while preserving the binaural cues of the target signal.
+We then extend the evaluation to reverberant conditions.
+Our method proves robust, maintaining speech clarity and source directionality while simultaneously reducing reverberation.
+A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods demonstrates that our approach attains performance on par with competing techniques in terms of noise reduction and perceptual quality, while offering a clear advantage in preserving binaural cues.
 Demo-page: https://bi-ctse-hrtf.github.io
@@
 While this ability may pose a challenge even to those with normal hearing, it presents a far greater obstacle for individuals with hearing impairments.
-Binaural Target Speaker Extraction (Bi-TSE) aims at selectively extracting one desired speaker of the multiple competing speakers, in a binaural setting.
+Binaural Target Speaker Extraction (Bi-TSE) aims to selectively extract the desired speaker from multiple competing speakers in a binaural setting.
 In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3].
-Most hearing aids feature multiple microphones on each earpiece and fall into one of the two categories: 1) Bilateral systems, in which each device operates independently, yielding a single-channel output.
-2) Shared systems, where both earpieces share all microphones, producing a dual-channel output.
+Most hearing aids employ either a single microphone or multiple microphones per earpiece and can be broadly categorized into two groups: 1) Bilateral systems, where each device operates independently and provides a single-channel output per earpiece; 2) Shared systems, where all microphones across both earpieces are jointly utilized, yielding a dual-channel output.
 Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears.
@@
 More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
-Unlike Blind Source Separation (BSS) algorithms, which separate all sources, Target Speaker Extraction (TSE) algorithms leverage prior knowledge of the target speaker(s) to isolate their voice.
-The prior knowledge about the speaker—commonly referred to as “clue” [6]—can include any additional information about the target speaker.
-For instance, it might be visual data depicting the speaker’s appearance as in [7] that uses a face image to extract its corrospondig speech, or [8] that uses both audio and visual clues, where the audio clue is an enrollment utterance and the visual clue is a video of the speaker speaking within the mixture.
-Another popular clue is a pre-recorded sample used to extract speaker-specific features, as in [9, 10, 11, 12].
-While examining prior works, such as [13], we observed the particular challenge of separating two voices with highly similar characteristics—for example, speakers of the same gender.
+Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge of the target speaker(s) to isolate their voice.
+This prior knowledge—commonly referred to as a “clue” [6]—can take various forms.
+A widely used clue is a pre-recorded speech sample, often referred to as an enrollment utterance, which provides speaker-specific features [7, 8, 9, 10].
+Clues may also come from other modalities, such as visual data.
+For example, [11] uses a face image to extract the corresponding speech, while [12] combines audio and visual clues, with the audio clue being an enrollment utterance and the visual clue a video of the speaker speaking within the mixture.
+While examining prior works, such as [13], we observed a particular challenge in separating two voices with highly similar characteristics, for example, speakers of the same gender.
 Consequently, we opted for an approach that does not depend on speaker–embedding–based clues.
 Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better.
-For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA.
+For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA and extracts spatial features from it.
 These spatially guided methods consistently outperform single-microphone solutions.
-In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that leverages the listener’s HRTF, aligned with the target speaker’s DOA, as an extraction clue.
+In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction cue.
 Our approach employs a complex‑valued DNN that operates directly in the STFT domain.
-Unlike real‑valued models that treat real and imaginary parts separately—or architectures that rely solely on the magnitude spectrogram—our network processes the full complex representation, more effectively preserving phase information.
 The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved.
 Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics.
-The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases.
+The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
 II Problem Formulation
@@
 Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation).
-Similarly, the interfering speaker’s DOA is ($\theta_{i}$,$\phi_{i}$).
+Similarly, the interfering speaker’s DOA is ($\theta_{i}$$\phi_{i}$).
 Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively.
-This signal can be written as
+The binaural signal can be written as
 $\bm{x}_{b}(n,k)\;=\;s_{d}(n,k)\,\bm{h}(\theta_{d},\phi_{d},k)+s_{i}(n,k)\,\bm{h}(\theta_{i},\phi_{i},k)$ (1)
-where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech of the desired speaker and the interfering speaker, respectively.
-Let $\gamma=\{d,i\}$ denote the speaker index, either the desired ($d$) or the interferer ($i$).
-Here, $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ is a Transfer Function (TrF) modeling acoustic propagation from source $\gamma$ to both earpieces as a superposition of HRTFs associated with the DOAs of all reflections.
-The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, represents the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$.
+where $s_{d}(n,k)$ and $s_{i}(n,k)$ are the anechoic speech signals of the desired speaker and the interfering speaker, respectively.
+Let $\gamma\in\{d,i\}$ denote the speaker index, where $d$ refers to the desired source and $i$ to the interferer.
+The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all reflections.
+The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizes the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$.
 According to the precedence effect [16], the direct-path HRTF dictates the perceived source DOA.
 Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity.
-Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$ Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
-Our work has three main objectives: 1) to extract the target speaker at a known DOA from binaural input using the listener’s HRTF; 2) to generate a binaural output that preserves the spatial cues of the desired source; and 3) to dereverberate the output, retaining only the direct-path HRTF.
+Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$.
+Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
+Given the HRTF corresponding to the desired source’s DOA $(\theta_{d},\phi_{d})$, namely the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$.
+The extracted signal is denoted by $\hat{\tilde{\bm{s}}}_{d}$.
+This binaural output is required to: 1) preserve the spatial cues of the desired source, and 2) dereverberate the output by retaining only the direct-path HRTF, thereby also dereverberating the desired speech signal.
 III Proposed Method
-We propose a deep neural network model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
-Given the HRTF from the desired source DOA $(\theta_{d},\phi_{d})$, namely $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$.
-We denote the extracted signal as $\hat{\tilde{\bm{s}}}_{d}$.
+We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
 III-A Architecture
-The proposed model architecture is based on the approach presented in [14], modified to fully handle complex-valued inputs, rather than the real and imaginary components independently.
+The proposed model architecture is based on the approach presented in [14], modified to fully handle dual-channel, complex-valued inputs, rather than the real and imaginary components independently.
+The proposed model employs a U-Net architecture, incorporating the HRTF at the bottleneck.
+The individual modules comprising the architecture are described in detail below.
 Encoder-Decoder Blocks of the U-net: The encoder comprises five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation function:
@@
 However, unlike traditional methods, we implement these self-attention layers directly on complex-valued tensors by replacing the softmax activation function with the complex-valued $\mathbb{C}\text{ReLU}$.
-To capture cross-channel (left and right) information in the HRTF, we apply the self-attention mechanism that enables the channels to attend to each other through learned attention weights.
+To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights.
 The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal.
@@
 Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage.
-A block diagram of this architecture is presented in Fig.
+A block diagram of the proposed architecture is presented in Fig.
 1.
 […]
-Figure 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed signal at the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of estimated desired signal.
+Fig.
+1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated binaural desired signal.
 III-B Loss function
 We initially trained the proposed model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
-[…]
+$\text{SI-SDR}(\bm{s},\hat{\bm{s}})=20\log_{10}\left(\frac{\|\beta\bm{s}\|}{\|\beta\bm{s}-\hat{\bm{s}}\|}\right);\quad\beta=\frac{\hat{\bm{s}}^{\top}\bm{s}}{\|\bm{s}\|^{2}}\,$, (3)
 where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all samples of the target and estimated signals, respectively.
@@
 $\mathcal{L}=\frac{1}{2}\sum_{\gamma=\{d,i\}}\left(-\text{SI-SDR}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})+\alpha\cdot\mathcal{L}_{\text{MAE}}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})\right).$ (5)
-Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a two-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$).
+Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$).
 Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality.
+The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
 IV Experimental Study
@@
 where $L$ is the reflection order. The direct path corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$.
-The RIEC HRTF Dataset [19, 20]: A measured HRTF database consists of 105 subjects, each with 865 directions.
+The RIEC HRTF Dataset [19, 20]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions.
 For this study, we focused on a single subject.
 The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
+We employed the SofaMyRoom framework [23] to simulate both anechoic and reverberant conditions, following the formulation in (6).
+This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time.
+It also supports zero-order simulations for anechoic scenarios.
+The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room.
+The overall simulation setup is illustrated in Fig.
+2.
+Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [24].
+To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
+IV-B Algorithm Setup
+The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, then transformed with a 512-sample STFT (75% overlap).
+The model is trained in a supervised manner: in the anechoic setting, targets are generated by convolving the desired speech with the corresponding HRTF; in the reverberant setting, by simulating a room and keeping only the first arrival (zero-order reflection).
+The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and converted to the frequency domain via Fast Fourier Transform (FFT).
 -90
@@
 Refer to captionRefer to captionRefer to caption
-Figure 2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$ [^2] Images sourced from: https://www.freepik.com
-We employed the SofaMyRoom framework [23] to simulate reverberant and anechoic conditions, similar to (6).
-This tool places a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a target reverberation time.
-It also supports zero-order simulations for anechoic scenarios.
-For training, we sampled $T_{60}\sim\mathcal{U}[0.2,0.8]\,$s, and for testing, $T_{60}\sim\mathcal{U}[0.1,0.4]\,$s.
-Speaker and listener positions were also drawn uniformly within the room’s dimensions.
-The simulation arrangement is illustrated in Fig.
-2.
-Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [24].
-To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
-IV-B Algorithm Setup
-The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, and transformed using a 512-sample STFT with 75% overlap.
-Our model was trained in a supervised manner.
-For the anechoic setting, the target signals were generated by convolving the desired speech signal with the corresponding HRTF.
-For the reverberant setting, the targets were obtained by simulating a reverberant room and using only the first arrival (i.e., zero-order reflection).
-The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and transformed to the frequency domain using Fast Fourier Transform (FFT).
+Fig.
+2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$.
+Images source: freepik.com
 IV-C Metrics
-To evaluate our method’s performance, we used SI-SDR[27], Perceptual Evaluation of Speech Quality (PESQ)[28], and Short-time Objective Intelligibility (STOI)[29] as measures of speech quality and intelligibility.
-lFurthermore, to verify that the estimated signals retain their binaural characteristics, we compare the Probability Density Function (p.d.f.) of the ILD and ITD of the target signals and the extracted signals.
-The similarity of the p.d.f.s indicates that the spatial cues are preserved.
+To evaluate our method, we employed SI-SDR [27], wide-band and narrow-band Perceptual Evaluation of Speech Quality (PESQ[28], and Short-time Objective Intelligibility (STOI[29] as intrusive metrics of speech quality and intelligibility, i.e., measures that compare the extracted signal against a reference target.
+In addition, we used DNS-MOS [30], a non-intrusive DNN-based predictor of Mean Opinion Score (MOS), which estimates speech quality (SIG), background noise quality (BAK), and overall quality (OVRL).
+[…]
+(a) Reverberant Mixture
+(b) $\tilde{\bm{s}}_{d}$
+(c) $\hat{\tilde{\bm{s}}}_{d}$
+(d) $\tilde{\bm{s}}_{i}$
+(e) $\hat{\tilde{\bm{s}}}_{i}$
+Ground-truthBi-TSE-HRTF (proposed)BDE-BiTSE[15]BLCMV Beamformer [4]
+Fig.
+3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced.
+$\tilde{\bm{s}}_{d}$ and $\tilde{\bm{s}}_{i}$ denotes the anechoic target signals.
+Graphs produced by [31].
+Table I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples.
+Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap .
+Best scores are highlighted in bold.
+[…]
+LS Unprocessed (2.688\3.380\3.140) -0.01 1.23 1.61 0.72 - -
+BLCMV Beamformer [4] (2.827\3.304\3.495) -1.0 3.46 3.9 0.95 0.672 0.599
+BDE-BiTSE [15] (3.072\3.471\3.768) 24.01 3.66 4.02 0.99 0.142 0.485
+Bi-TSE-HRTF (proposed) (2.999\3.398\3.742) 20.18 3.14 3.78 0.98 0.1024 0.0042
+MLS-French Unprocessed (2.499\3.252\2.947) -0.01 1.23 1.63 0.71 - -
+BLCMV Beamformer (2.761\3.277\3.384) -1.03 3.37 3.82 0.94 0.694 0.608
+BDE-BiTSE (2.987\3.403\3.709) 23.67 3.67 4.04 0.99 0.3 0.503
+Bi-TSE-HRTF (proposed) (2.900\3.324\3.655) 19.86 3.14 3.79 0.98 0.1039 0.0041
+WSJ0 Unprocessed (2.845\3.376\3.411) 0 1.25 1.57 0.72 - -
+BLCMV Beamformer (3.044\3.444\3.748) -0.42 3.48 3.84 0.95 0.745 0.613
+BDE-BiTSE (3.230\3.543\3.951) 25.59 3.77 4.01 1.0 0.126 0.484
+Bi-TSE-DOA (2.965\3.376\3.660) 16.44 2.43 3.09 0.97 0.165 0.990
+Bi-TSE-HRTF-RI (3.110\3.504\3.764) 21.39 2.93 3.57 0.99 0.085 0.0044
+Bi-TSE-HRTF (proposed) (3.18\3.501\3.911) 21.41 3.47 3.88 0.99 0.0035 0.0756
+WSJ0 (Reverb) Unprocessed (2.348\3.069\2.783) -1.9 1.15 1.44 0.69 - -
+BLCMV Beamformer (1.840\2.612\2.038) -1.11 1.35 1.88 0.74 0.430 0.593
+BDE-BiTSE (3.141\3.438\3.952) 14.8 3.09 3.56 0.98 0.204 0.386
+Bi-TSE-DOA (2.681\3.196\3.324) 8.29 1.82 2.44 0.91 0.296 0.990
+Bi-TSE-HRTF-RI (3.030\3.362\3.838) 11.18 2.01 2.74 0.95 0.174 1.001
+Bi-TSE-HRTF (proposed) (2.86\3.348\3.481) 10.67 2.17 2.79 0.94 0.186 0.0082
 IV-D Competing Methods
-To ensure a fair comparison, all baseline methods operate under an identical setup, namely, binaural TSE using only two microphones (one per ear).
-Accordingly, we compare our method against the BLCMV beamformer [4] and a recent DNN-based model [30], both designed specifically for this setup.
+To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear).
+Accordingly, we compared our method against the BLCMV beamformer [4] and a recent DNN-based model [15], both specifically designed for this scenario.
+The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the same training data as ours.
+We ensured that the strongest variant reported in the paper (BDE+CDF+IPD+SDF) was used for comparison.
 In addition, we trained two variants of our model.
-The first is Bi-TSE-HRTF-RI, a real-valued model, in which the complex-value input features are split into Real-Imaginary (RI) components and concatenated, resulting in four input channels instead of the two input channels in the original scheme.
-The same procedure is applied to the HRTF clue.
-The second is Bi-TSE-DOA, a complex-valued model that incorporates the DOA as a clue instead of the HRTF.
-Here, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs and passed through a fully-connected layer to match the dimensionality of the bottleneck.
-For the specified variants, both terms of (5) remain intact, i.e., the target signal is the binaural rendering of the desired speech signal with the correct HRTF.
-(a) SI-SDR[dB]
-(b) PESQ
-(c) STOI
-Figure 3: Histograms of speech quality and intelligibility measures, before and after the application of the proposed TSE algorithm for the anechoic setting.
-TABLE I: The evaluation datasets include 3k speech samples mixed with SIR sampled from $\mathcal{U}[0,5]$ dB.
-Speaker overlap is either sampled from $\mathcal{U}[0\%,100\%]$ (designated by) or set to 100% (designated by∗∗), with audio lengths of 5 s and 8 s, respectively.
-Best scores per metric and dataset are highlighted in bold.
-Method Dataset Reverberant SI-SDR (dB) $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ $\Delta$ILD [dB] $\downarrow$ $\Delta$ITD [mS] $\downarrow$
-Unprocessed LS∗∗ 2.4 1.36 0.78 - -
-BLCMV Beamformer [4] LS∗∗ 0.86 3.93 0.95 0.7341 0.6032
-Bi-CSim-TSE [30] LS 18.52 3.05 0.96 - -
-Bi-TSE-HRTF (proposed) LS∗∗ 19.35 3.27 0.98 0.1311 0.0033
-Unprocessed MLS-French∗∗ 2.37 1.29 0.77 - -
-BLCMV Beamformer [4] MLS-French∗∗ 0.98 3.92 0.95 0.7092 0.5829
-Bi-TSE-HRTF (proposed) MLS-French∗∗ 18.9 3.23 0.98 0.2697 0.0120
-Unprocessed WSJ0∗∗ 0.06 1.23 0.73 - -
-BLCMV Beamformer [4] WSJ0∗∗ 1.51 3.98 0.96 0.733 0.5669
-Bi-TSE-DOA WSJ0∗∗ 16.3 2.47 0.97 0.1304 1.0145
-Bi-TSE-HRTF-RI WSJ0∗∗ 21.36 2.97 0.99 0.0934 0.00435
-Bi-TSE-HRTF (proposed) WSJ0∗∗ 20.37 3.49 0.99 0.1003 0.0027
-Unprocessed WSJ0∗∗ -1.15 1.21 0.7 - -
-BLCMV Beamformer [4] WSJ0∗∗ 1.37 1.6 0.82 0.3303 0.5439
-Bi-TSE-DOA WSJ0∗∗ 9.53 2.08 0.93 0.2410 1.0125
-Bi-TSE-HRTF (proposed) WSJ0∗∗ 10.37 2.60 0.95 0.1998 0.0036
+The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are split into their RI components and concatenated, resulting in four input channels instead of two.
+The same decomposition is applied to the HRTF clue.
+The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue instead of the HRTF.
+In this case, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs, and passed through a fully connected layer to match the dimensionality of the bottleneck.
+For both variants, the loss function in (5) remains unchanged, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
 IV-E Results
-In Fig.
-3, we show histograms of the SI-SDR, PESQ, and STOI metrics before and after applying our TSE algorithm in the anechoic scenario.
-All three distributions shift markedly to the right, demonstrating substantial gains across signal quality, perceptual quality, and intelligibility.
-This confirms the algorithm’s effectiveness in isolating the desired speaker.
-In Table I, we compare our method against competing algorithms across all datasets under both anechoic and reverberant conditions.
-Our approach achieves significant gains on every metric except PESQ, where the BLCMV beamformer slightly outperforms the proposed method in anechoic case.
-This may be attributed to its distortionless response design criterion.
-When comparing the two additional variants of our model, denoted as Bi-TSE-DOA and Bi-TSE-HRTF-RI in Table I, we observe that, in contrast to the HRTF-guided model, the DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and is unable to preserve the binaural cues.
-We evaluated this model under reverberant conditions to emphasize that the HRTF is intrinsically tied to our problem formulation.
-The superior performance of the HRTF-guided network supports these findings.
-On the other end, the RI model achieves comparable SI-SDR and binaural cue preservation to the complex-valued model but suffers from lower output quality, as shown by its lower PESQ score in the anechoic setting.
-Remarkably, our approach demonstrates cross-dataset generalizability: although our model was trained exclusively on WSJ0-CSR1, it performs exceptionally well on other corpora, including a French dataset.
-We attribute this generalization to the absence of speaker-specific priors in our approach.
-Figure 4: A polar plot showing the SI-SDR of the output signals for each DOA, with the desired speaker located at $\theta_{d}=-75^{\circ}$ (red) and the interfering speaker at $\theta_{i}=-10^{\circ}$ (blue).
-The bold circle represents the 0 dB SI-SDR threshold.
-When inspecting the extracted signals, we found that the model effectively enhances the target speaker while nearly eliminating the interferer.
-To quantify its spatial resolution, we devised a “scanning” procedure: we iteratively extract signals using HRTFs corresponding to azimuth angles in $[-90^{\circ},90^{\circ}]$ at a fixed elevation angle and compute the SI-SDR of each extraction against the true target.
-Plotting these values yields a beam-pattern-like diagram (Fig.
-4), which shows a pronounced peak at the correct DOA.
-Additionally, the width of the main lobe indicates that the model remains effective even when the HRTF clue deviates slightly from the true HRTF.
-The proposed method’s ability to preserve binaural cues is also evaluated and compared with the BLCMV, where applicable.
-Table I presents the mean ILD and ITD deviations.
-Figure 5 depicts the p.d.f.s of both cues for the target and extracted signals, showing a close match and confirming faithful spatial cue preservation.
-(a) Mixture
-(b) $\hat{\tilde{\bm{s}}}_{d}$
-(c) $\hat{\tilde{\bm{s}}}_{i}$
-Figure 5: The p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=90^{\circ}$, $\theta_{i}=15^{\circ}$, and $\phi_{d,i}=0^{\circ}$. Graphs produced by [31].
+In Table I, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions.
+In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE.
+Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
+Remarkably, both DNN-based methods demonstrate strong cross-dataset generalizability.
+Although the models were trained exclusively on WSJ0-CSR1, they perform well on other corpora, including a French dataset.
+We attribute this generalization to the absence of speaker-specific priors and to the reliance on a known HRTF database.
+[…]
+[…]
+BLCMV Beamformer 1.7 >90>90 >150>150
+BDE-BiTSE 1.0667 30 90
+Bi-TSE-HRTF 0.0 -60 0
+Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI.
+The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation.
+By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation.
+Figure 3 shows the joint ITDILD p.d.f.s across methods.
+The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues.
+In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations.
+The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f.
+becomes more concentrated, indicating dominance of the direct path in the output of the proposed method.
+Overall, it is evident that the proposed approach achieves this more effectively than competing techniques.
+Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig.
+3.
+Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec.
+IV-C.
+The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
+Fig.
+4: A polar plot of SI-SDR versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue).
+Stars mark true DOAs; the dotted arc denotes the 0,dB threshold.
+Finally, to quantify the spatial resolution of our model, we employ a “scanning” procedure: signals are iteratively extracted using HRTFs at azimuths within $90^{\circ},90^{\circ}]$ (with fixed elevation), and the SI-SDR of each extraction is computed against the target.
+Plotting these values produces a beam-pattern-like diagram (Fig.
+4) that exhibits a clear peak at the true DOA.
+The width of the main lobe further characterizes the model’s robustness when the HRTF cue deviates slightly from the ground-truth.
 V Conclusions
-This paper presents a speaker‑independent TSE method that leverages the listener’s HRTFs in both anechoic and reverberant settings, resulting in a personalized solution for the proposed problem.
-Through cross‑dataset evaluations, we demonstrate its robust generalization across diverse corpora.
-Our approach not only delivers high‑quality speech but also faithfully preserves the binaural character of the signal.
-Comparisons with other binaural TSE techniques underscore the importance of spatial cues tied to the listener’s HRTF for accurate and consistent target‑speaker extraction under varying acoustic conditions.
+This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, yielding a personalized solution to the task at hand.
+Through cross-dataset evaluations, we demonstrate robust generalization across diverse corpora.
+The proposed approach not only produces high-quality speech but also faithfully preserves the binaural characteristics of the signal.
+Comparative analyses with existing binaural TSE methods reveal substantial improvements in spatial consistency, while preserving high speech quality.
+These findings highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction under a wide range of acoustic conditions.

### v3

@@
 Demo-page: https://bi-ctse-hrtf.github.io
-Index Terms:
-Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
-I Introduction
+Index Terms—  Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
+1 Introduction
 Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking.
@@
 The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
-II Problem Formulation
+2 Problem Formulation
 We address the problem of multiple concurrent speakers, captured by two microphones, each mounted on one of the listener’s ears in a noiseless environment.
@@
 This binaural output is required to: 1) preserve the spatial cues of the desired source, and 2) dereverberate the output by retaining only the direct-path HRTF, thereby also dereverberating the desired speech signal.
-III Proposed Method
+3 Proposed Method
 We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
-III-A Architecture
+3.1 Architecture
 The proposed model architecture is based on the approach presented in [14], modified to fully handle dual-channel, complex-valued inputs, rather than the real and imaginary components independently.
@@
 We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal.
-Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage.
-A block diagram of the proposed architecture is presented in Fig.
+Finally, the resulting tensor undergoes four consecutive self-attention layers and a fully connected layer, preparing it for the decoding stage.
+A block diagram of the proposed architecture appears in Fig.
 1.
@@
 Fig.
-1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated binaural desired signal.
-III-B Loss function
-We initially trained the proposed model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
+1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated desired signal.
+3.2 Loss function
+We initially trained the model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
 $\text{SI-SDR}(\bm{s},\hat{\bm{s}})=20\log_{10}\left(\frac{\|\beta\bm{s}\|}{\|\beta\bm{s}-\hat{\bm{s}}\|}\right);\quad\beta=\frac{\hat{\bm{s}}^{\top}\bm{s}}{\|\bm{s}\|^{2}}\,$, (3)
@@
 The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
-IV Experimental Study
+4 Experimental Study
 In this section, we describe the data-generation process and experimental setup, evaluate the performance of the proposed approach, and compare it with previous TSE methods.
-IV-A Data Generation
+4.1 Data Generation
 Binaural Impulse Rsponses: The binaural Impulse Responses are formed by superimposing attenuated and time-shifted HRIRs, corresponding to acoustic reflections generated via the image method [18], as expressed in (6).
@@
 To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
-IV-B Algorithm Setup
-The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, then transformed with a 512-sample STFT (75% overlap).
-The model is trained in a supervised manner: in the anechoic setting, targets are generated by convolving the desired speech with the corresponding HRTF; in the reverberant setting, by simulating a room and keeping only the first arrival (zero-order reflection).
-The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and converted to the frequency domain via Fast Fourier Transform (FFT).
 -90
@@
 Images source: freepik.com
-IV-C Metrics
+4.2 Algorithm Setup
+The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, then transformed with a 512-sample STFT (75% overlap).
+The model is trained in a supervised manner: in the anechoic setting, targets are generated by convolving the desired speech with the corresponding HRTF; in the reverberant setting, by simulating a room and keeping only the first arrival (zero-order reflection).
+The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and converted to the frequency domain via Fast Fourier Transform (FFT).
+4.3 Metrics
 To evaluate our method, we employed SI-SDR [27], wide-band and narrow-band Perceptual Evaluation of Speech Quality (PESQ[28], and Short-time Objective Intelligibility (STOI[29] as intrusive metrics of speech quality and intelligibility, i.e., measures that compare the extracted signal against a reference target.
@@
 Graphs produced by [31].
-Table I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples.
+Table 1: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples.
 Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap .
@@
 Bi-TSE-HRTF (proposed) (2.86\3.348\3.481) 10.67 2.17 2.79 0.94 0.186 0.0082
-IV-D Competing Methods
+s
+[…]
+[…]
+BLCMV Beamformer 1.7 >90>90 >150>150
+BDE-BiTSE 1.0667 30 90
+Bi-TSE-HRTF 0.0 -60 0
+4.4 Competing Methods
 To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear).
@@
 For both variants, the loss function in (5) remains unchanged, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
-IV-E Results
-In Table I, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions.
+4.5 Results
+In Table 1, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions.
 In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE.
 Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
-Remarkably, both DNN-based methods demonstrate strong cross-dataset generalizability.
-Although the models were trained exclusively on WSJ0-CSR1, they perform well on other corpora, including a French dataset.
-We attribute this generalization to the absence of speaker-specific priors and to the reliance on a known HRTF database.
-[…]
-[…]
-BLCMV Beamformer 1.7 >90>90 >150>150
-BDE-BiTSE 1.0667 30 90
-Bi-TSE-HRTF 0.0 -60 0
+Both DNN-based methods show strong cross-dataset generalization: though trained only on WSJ0-CSR1, they perform well on other corpora, including French, as shown in Table 1.
+We attribute this to the absence of speaker-specific priors and the reliance on a known HRTF database.
 Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI.
@@
 Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec.
-IV-C.
+4.3.
 The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
@@
 Stars mark true DOAs; the dotted arc denotes the 0,dB threshold.
-Finally, to quantify the spatial resolution of our model, we employ a “scanning” procedure: signals are iteratively extracted using HRTFs at azimuths within $90^{\circ},90^{\circ}]$ (with fixed elevation), and the SI-SDR of each extraction is computed against the target.
-Plotting these values produces a beam-pattern-like diagram (Fig.
-4) that exhibits a clear peak at the true DOA.
-The width of the main lobe further characterizes the model’s robustness when the HRTF cue deviates slightly from the ground-truth.
-V Conclusions
+Finally, to quantify spatial resolution, we use a “scanning” procedure: signals are extracted with HRTFs at azimuths in $90^{\circ},90^{\circ}]$, and the SI-SDR of each is computed against the target.
+Plotting these values yields a beam-pattern-like diagram (Fig.
+4) with a clear peak at the true DOA.
+The main lobe width characterizes the model’s stability when the HRTF clue deviates mildly from the ground truth.
+5 Conclusions
 This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, yielding a personalized solution to the task at hand.

### v4

@@
-Binaural Target Speaker Extraction using HRTFs
+Yoav Ellinson    Sharon Gannot
 Abstract
-In this work, we aim to imitate the human ability to selectively attend to a single speaker, even in the presence of multiple simultaneous talkers.
-To achieve this, we propose a novel approach for binaural target speaker extraction that leverages the listener’s Head-Related Transfer Function (HRTF) to isolate the desired speaker.
-Notably, our method does not rely on speaker embeddings, making it speaker-independent and enabling strong generalization across multiple speech datasets and languages.
+In this work, we address the problem of binaural target-speaker extraction in the presence of multiple simultaneous talkers.
+We propose a novel approach that leverages the individual listener’s Head-Related Transfer Function (HRTF) to isolate the target speaker.
+The proposed method is speaker-independent, as it does not rely on speaker embeddings.
 We employ a fully complex-valued neural network that operates directly on the complex-valued Short-Time Fourier transform (STFT) of the mixed audio signals, and compare it to a Real-Imaginary (RI)-based neural network, demonstrating the advantages of the former.
@@
 Our method proves robust, maintaining speech clarity and source directionality while simultaneously reducing reverberation.
-A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods demonstrates that our approach attains performance on par with competing techniques in terms of noise reduction and perceptual quality, while offering a clear advantage in preserving binaural cues.
+A comparative analysis with existing binaural Target Speaker Extraction (TSE) methods shows that the proposed approach achieves performance comparable to state-of-the-art techniques in terms of noise reduction and perceptual quality, while providing a clear advantage in preserving binaural cues.
 Demo-page: https://bi-ctse-hrtf.github.io
-Index Terms—  Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
-1 Introduction
-Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments where multiple speakers are actively speaking.
+I Introduction
+Humans have a remarkable ability to focus on the voice of a single speaker, even in complex auditory environments with multiple speakers.
 Such a scenario is often referred to as the cocktail party problem [1].
@@
 Binaural Target Speaker Extraction (Bi-TSE) aims to selectively extract the desired speaker from multiple competing speakers in a binaural setting.
-In this context, preserving the binaural cues is critical: hearing aid and hearing aid algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3].
+In this context, preserving the binaural cues is critical: hearing aids and their associated algorithms must maintain the listener’s spatial awareness to ensure natural sound perception [2, 3].
 Most hearing aids employ either a single microphone or multiple microphones per earpiece and can be broadly categorized into two groups: 1) Bilateral systems, where each device operates independently and provides a single-channel output per earpiece; 2) Shared systems, where all microphones across both earpieces are jointly utilized, yielding a dual-channel output.
 Two important spatial cues are the Interaural Level Difference (ILD) and Interaural Time Difference (ITD), which capture magnitude and arrival-time differences between the left and right ears.
-In bilateral systems, where each device processes its own channel independently, it is particularly challenging to preserve these binaural cues.
+In bilateral systems, in which each device processes its own channel independently, it is particularly challenging to preserve binaural cues.
 In contrast, Multi input Multi output (MIMO) solutions jointly process both channels, yielding more natural and spatially accurate outputs.
@@
 More recent work, such as [5], employs a Deep Neural Network (DNN) to process all microphone signals and produce left- and right-channel outputs, effectively preserving spatial information.
-Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge of the target speaker(s) to isolate their voice.
+Unlike Blind Source Separation (BSS) algorithms, which separate all sources, TSE algorithms leverage prior knowledge on the target speaker(s) to isolate their voice.
 This prior knowledge—commonly referred to as a “clue” [6]—can take various forms.
@@
 While examining prior works, such as [13], we observed a particular challenge in separating two voices with highly similar characteristics, for example, speakers of the same gender.
-Consequently, we opted for an approach that does not depend on speaker–embedding–based clues.
-Extending beyond a single microphone enables us to exploit spatial information and isolate the desired speaker better.
-For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15] relies on the DOA and extracts spatial features from it.
+Consequently, we opted for an approach that does not rely on speaker–embedding–based clues.
+Extending beyond a single microphone enables us to exploit spatial information and more effectively isolate the desired speaker.
+For instance, [14] uses either the Relative Transfer Function (RTF) or the Direction of Arrival (DOA) of the target speaker relative to a microphone array, while [15, 16] rely on the DOA of the desired speaker and use it as a spatial clue.
 These spatially guided methods consistently outperform single-microphone solutions.
-In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction cue.
+In this work, we propose Binaural Target Speaker Extraction using HRTFs (Bi-TSE-HRTF), a target speaker extraction method that exploits the individual listener’s HRTF, aligned with the target speaker’s DOA, as the primary extraction clue.
 Our approach employs a complex‑valued DNN that operates directly in the STFT domain.
 The proposed architecture produces binaural audio outputs, ensuring that spatial cues are fully preserved.
-Moreover, in reverberant environments, the model isolates the direct-path’s HRTF, effectively dereverberating the signal and further enhancing its binaural characteristics.
+Moreover, in reverberant environments, the model isolates the direct path’s HRTF, thereby dereverberating the signal and hence further enhancing its binaural characteristics.
 The method is thoroughly assessed using simulated Room Impulse Responses and standard HRTF databases, and compared with two competing methods, a beamformer-based solution and a DNN-based solution.
-2 Problem Formulation
-We address the problem of multiple concurrent speakers, captured by two microphones, each mounted on one of the listener’s ears in a noiseless environment.
-For clarity, we focus on the simpler case of two concurrent speakers, as formulated below.
-Denote the desired speaker’s DOA by $\theta_{d}$ (azimuth) and $\phi_{d}$ (elevation).
-Similarly, the interfering speaker’s DOA is ($\theta_{i}$$\phi_{i}$).
-Let $\bm{x}_{b}(n,k)$ be the STFT of the binaural signal, with $n=0,1,\ldots,N-1$ and $k=0,1,\ldots,K-1$ the time-frame and frequency-bin indexes, respectively.
-The binaural signal can be written as
+II Problem Formulation
+We address the problem of multiple concurrent speakers captured by two microphones mounted on the listener’s ears, assuming a noiseless environment and a fixed source–listener radius.
+For clarity, we focus on the case of two concurrent speakers, as formulated below.
+Denote the desired speaker’s DOA by $(\theta_{d},\phi_{d})$ and the interferer’s by $(\theta_{i},\phi_{i})$.
+Let $\bm{x}_{b}(n,k)$ denote the STFT of the binaural signal, with $n\in[0,N-1]$ and $k\in[0,K-1]$ the time-frame and frequency-bin indices, respectively.
+The binaural signal is therefore formulated as:
 $\bm{x}_{b}(n,k)\;=\;s_{d}(n,k)\,\bm{h}(\theta_{d},\phi_{d},k)+s_{i}(n,k)\,\bm{h}(\theta_{i},\phi_{i},k)$ (1)
@@
 Let $\gamma\in\{d,i\}$ denote the speaker index, where $d$ refers to the desired source and $i$ to the interferer.
-The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all reflections.
-The HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizes the listener’s frequency response to a wave arriving from DOA $(\theta,\phi)$.
+The acoustic transfer from source $\gamma$ to both earpieces is described by $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, which models the Transfer Function (TrF) as a superposition of HRTFs corresponding to the DOAs of all acoustic reflections in the enclosure, with the HRTF, $\bm{h}_{\text{hrtf}}(\theta,\phi,k)$, characterizeing the listener’s response to a wave arriving from DOA $(\theta,\phi)$.
 According to the precedence effect [17], the direct-path HRTF dictates the perceived source DOA.
 Thus, although $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$ encompasses all reflections, it is indexed by the direct-path DOA for clarity.
-Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$.
+Finally, define $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)$, with $t$ the filter’s tap index, as the time-domain representation of $\bm{h}(\theta_{\gamma},\phi_{\gamma},k)$, and similarily define the Head-Related Impulse Response (HRIR) at the DOA $\theta,\phi$ as $\bar{\bm{h}}_{\text{hrir}}(\theta,\phi,t)$.
 Finally, the target signal, $\tilde{\bm{s}}_{d}=[\tilde{s}_{d}^{l},\tilde{s}_{d}^{r}]^{\top}$, is a concatenation of the desired speech signal as received at the left and right ears, respectively.
-Given the HRTF corresponding to the desired source’s DOA $(\theta_{d},\phi_{d})$, namely the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$.
+Given the individual HRTF of the hearing-aid wearer corresponding to the desired source DOA $(\theta_{d},\phi_{d})$, denoted by the binaural response $\bm{h}_{\text{hrtf}}(\theta_{d},\phi_{d},k)$, our method aims to extract $\tilde{\bm{s}}_{d}$ from the mixed signal $\bm{x}_{b}$.
 The extracted signal is denoted by $\hat{\tilde{\bm{s}}}_{d}$.
-This binaural output is required to: 1) preserve the spatial cues of the desired source, and 2) dereverberate the output by retaining only the direct-path HRTF, thereby also dereverberating the desired speech signal.
-3 Proposed Method
-We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue, effectively leveraging spatial information for the extraction task.
-3.1 Architecture
-The proposed model architecture is based on the approach presented in [14], modified to fully handle dual-channel, complex-valued inputs, rather than the real and imaginary components independently.
+It is required that the binaural output satisfy the following conditions: 1) preservation of the spatial cues of the desired source, and 2) dereverberation of the desired speech output by retaining only the direct-path HRTF.
+In this work, we focus on individualized HRTF database, namely the set of HRTFs of the hearing aid wearer across all azimuth and elevation directions, which should be given to the algorithm.
+HRTF ersonalization using anthropometric measurements [18].
+III Proposed Method
+We propose a DNN-based model to extract a desired speech signal corresponding to a given HRTF clue.
+III-A Architecture
+The proposed model architecture is built upon the approach presented in [14], with modifications to handle dual-channel, complex-valued inputs rather than treating the real and imaginary components independently.
 The proposed model employs a U-Net architecture, incorporating the HRTF at the bottleneck.
-The individual modules comprising the architecture are described in detail below.
-Encoder-Decoder Blocks of the U-net: The encoder comprises five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation function:
-$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\left(\mathcal{R}(z)\right)+\mathrm{j}\cdot\text{ReLU}\left(\mathcal{I}(z)\right).$ (2)
-Here $\text{ReLU}(x)=\max(0,x)$ and $\mathcal{R}(z),\mathcal{I}(z)$ denote the real and imaginary parts of z, respectively.
-Additionally, $\mathrm{j}$ represents the imaginary unit.
-Similarly, the decoder comprises five up-convolution blocks, each employing transposed convolutional layers.
-Skip connections are utilized between corresponding encoder and decoder blocks of the U-net structure to improve reconstruction accuracy.
-Additionally, the decoder includes a final convolutional layer with no activation function, enabling the representation of both positive and negative values at the output.
-This design choice is essential since the STFT representation inherently includes both positive and negative real and imaginary components.
-Bottelneck Processing: The extraction clue is applied to the encoded input features at the bottleneck, with self-attention layers employed both before and after its application.
-However, unlike traditional methods, we implement these self-attention layers directly on complex-valued tensors by replacing the softmax activation function with the complex-valued $\mathbb{C}\text{ReLU}$.
-To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights.
-The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal.
-The output of this process serves as the HRTF’s embedding.
+The individual modules comprising the architecture are described in detail below and depicted in Fig.
+1.
+Encoder–Decoder Blocks of the U-Net: The encoder consists of five down-convolution blocks, each comprising a 2D convolutional layer, global layer normalization, and a complex-valued ReLU activation defined as
+$\mathbb{C}\text{ReLU}(z)=\text{ReLU}\!\left(\mathcal{R}(z)\right)+\mathrm{j}\,\text{ReLU}\!\left(\mathcal{I}(z)\right),$ (2)
+where $\text{ReLU}(x)=\max(0,x)$, $\mathcal{R}(z)$ and $\mathcal{I}(z)$ denote the real and imaginary parts of $z$, and $\mathrm{j}$ is the imaginary unit.
+The decoder mirrors the encoder with five up-convolution blocks based on transposed convolutions.
+Skip connections are employed between corresponding encoder and decoder blocks to improve reconstruction accuracy.
+The decoder further includes a final convolutional layer without an activation function, allowing the output to represent both positive and negative real and imaginary values, which is required for the complex STFT representation.
+Bottelneck: The extraction clue is applied at the bottleneck of the encoded input features, with self-attention layers used both before and after its application.
+Unlike conventional approaches, these self-attention layers operate directly on complex-valued tensors by replacing the softmax activation with the complex-valued $\mathbb{C}\text{ReLU}$.
+To capture cross-channel (left–right) information in the HRTF, a self-attention mechanism is employed, enabling the channels to attend to one another via learned weights.
+The resulting matrix is flattened and projected via a fully-connected layer to match the dimensionality of the encoded mixture, yielding the HRTF embedding.
 Similarly, self-attention is applied to the encoded mixed signal.
-We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal.
-Finally, the resulting tensor undergoes four consecutive self-attention layers and a fully connected layer, preparing it for the decoding stage.
-A block diagram of the proposed architecture appears in Fig.
-1.
+The replicated HRTF embedding is then combined with each time frame of the encoded mixture through element-wise multiplication, thereby imposing the extraction clue.
+The resulting tensor is processed by four additional self-attention layers and a fully connected layer before decoding.
 […]
-Fig.
-1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated desired signal.
-3.2 Loss function
-We initially trained the model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
+Figure 1: A block diagram of the proposed method, where $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the HRTF of the desired speaker’s DOA, $\bm{x}_{b}$ represents the mixed binaural signal, and $\hat{\tilde{\bm{s}}}_{d}$ represents the estimated desired signal, both in the STFT domain.
+III-B Loss function
+The model was initially trained using only the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss (3).
 $\text{SI-SDR}(\bm{s},\hat{\bm{s}})=20\log_{10}\left(\frac{\|\beta\bm{s}\|}{\|\beta\bm{s}-\hat{\bm{s}}\|}\right);\quad\beta=\frac{\hat{\bm{s}}^{\top}\bm{s}}{\|\bm{s}\|^{2}}\,$, (3)
-where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all samples of the target and estimated signals, respectively.
+where $\bm{s}$ and $\hat{\bm{s}}$ are vectors comprising all time-domain samples of the target and estimated signals, respectively.
 However, relying solely on the SI-SDR loss proved challenging in preserving accurate spectral details, resulting in noticeable musical noise artifacts in the extracted signals.
@@
 $\mathcal{L}_{\text{MAE}}=\frac{1}{2KN}\sum_{c=\{l,r\}}\sum_{k=0}^{K-1}\sum_{n=0}^{N-1}\left||\tilde{s}_{d}^{c}(n,k)|-|\hat{\tilde{s}}_{d}^{c}(n,k)|\right|.$ (4)
-Here, $c=\{l,r\}$ denotes the channels, left or right.
+Here, $c=\{l,r\}$ denotes the output channels, left or right.
 To improve data utilization, both speakers were extracted from each mixed signal in each training step by alternating the HRTF clues and averaging the SI-SDR losses.
@@
 $\mathcal{L}=\frac{1}{2}\sum_{\gamma=\{d,i\}}\left(-\text{SI-SDR}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})+\alpha\cdot\mathcal{L}_{\text{MAE}}(\tilde{\bm{s}}_{\gamma},\hat{\tilde{\bm{s}}}_{\gamma})\right).$ (5)
-Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, where we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$).
+Due to the complexity of multi-objective optimization in high-dimensional space, we adopted a multi-stage training strategy, in which we first train the algorithm using only the SI-SDR loss (i.e., $\alpha=0$).
 Once a sufficient performance level has been reached, the $L_{1}$-MAE loss is introduced (namely, $\alpha>0$), improving both spectral accuracy and audio quality.
 The final stage fine-tunes the model using only the SI-SDR loss, as the $L_{1}$-MAE loss is prone to overfitting to the training data.
-4 Experimental Study
+IV Experimental Study
 In this section, we describe the data-generation process and experimental setup, evaluate the performance of the proposed approach, and compare it with previous TSE methods.
-4.1 Data Generation
+IV-A Data Generation
 Binaural Impulse Rsponses: The binaural Impulse Responses are formed by superimposing attenuated and time-shifted HRIRs, corresponding to acoustic reflections generated via the image method [20], as expressed in (6).
 $\bar{\bm{h}}(\theta_{\gamma},\phi_{\gamma},t)=\sum_{\ell=0}^{L}a_{\ell}\cdot\bar{\bm{h}}_{\text{hrir}}(\theta_{\ell},\phi_{\ell},t-\tau_{\ell})$ (6)
-where $L$ is the reflection order. The direct path corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$.
+where $L$ is the reflection order.
+The direct path ($\ell=0$) corresponds to the HRIR of the desired or interfering source at the listener’s head for the DOAs $(\theta_{\gamma},\phi_{\gamma})$, and the gain $a_{\ell}$ and the time delay $\tau_{\ell}$, obtained from the image method, are applied by scaling and time-shifting each HRTF component.
 The RIEC HRTF Dataset [21, 22]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions.
@@
 It also supports zero-order simulations for anechoic scenarios.
-The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room.
+The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and the listener position is drawn uniformly within the room.
 The overall simulation setup is illustrated in Fig.
@@
 Speech Datasets We trained and validated our model on the WSJ0‑CSR1 dataset [26].
-To evaluate its performance and generalizability, we tested on three additional corpora: 1) the test set of the LibriSpeech (LS) corpus [25], 2) the French subset of the Multilingual LibriSpeech (MLS) corpus [26], and 3) the WSJ0‑CSR1 test set.
 -90
@@
 Refer to captionRefer to captionRefer to caption
-Fig.
-2: A simulation setup example, illustrating two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$.
+Figure 2: A simulation illustration: two concurrent speakers: the desired speaker at $\theta_{d}=40^{\circ}$ (left) and the interferer at $\theta_{i}=-30^{\circ}$ (right), both at elevation $\phi$.
 Images source: freepik.com
-4.2 Algorithm Setup
+IV-B Algorithm Setup
 The speech signals, sampled at 16 kHz, are randomly selected, cropped, or zero-padded to 5 s, then transformed with a 512-sample STFT (75% overlap).
@@
 The HRTFs, used as the extraction clue, are downsampled from 48 kHz to 16 kHz and converted to the frequency domain via Fast Fourier Transform (FFT).
-4.3 Metrics
-To evaluate our method, we employed SI-SDR [27], wide-band and narrow-band Perceptual Evaluation of Speech Quality (PESQ[28], and Short-time Objective Intelligibility (STOI[29] as intrusive metrics of speech quality and intelligibility, i.e., measures that compare the extracted signal against a reference target.
-In addition, we used DNS-MOS [30], a non-intrusive DNN-based predictor of Mean Opinion Score (MOS), which estimates speech quality (SIG), background noise quality (BAK), and overall quality (OVRL).
-[…]
 (a) Reverberant Mixture
@@
 Ground-truthBi-TSE-HRTF (proposed)BDE-BiTSE[15]BLCMV Beamformer [4]
-Fig.
-3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced.
+Figure 3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced.
 $\tilde{\bm{s}}_{d}$ and $\tilde{\bm{s}}_{i}$ denotes the anechoic target signals.
 Graphs produced by [27].
-Table 1: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples.
+IV-C Metrics
+To evaluate our method, we use SI-SDR [28], wide- and narrow-band Perceptual Evaluation of Speech Quality (PESQ[29], and Short-time Objective Intelligibility (STOI[30] as intrusive metrics of speech quality and intelligibility.
+In addition, we employ DNS-MOS [31], a non-intrusive DNN-based Mean Opinion Score (MOS) predictor estimating speech quality (SIG), background noise quality (BAK), and overall quality (OVRL).
+Furthermore, to assess binaural cue preservation, we compare the ILDITD p.d.f.s of the extracted and target signals, where close agreement indicates preserved spatial cues.
+These metrics are computed using the procedure in [27].
+TABLE I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples.
 Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap.
-Best scores are highlighted in bold.
+Best scores are highlighted in boldface.
 […]
-LS Unprocessed (2.688\3.380\3.140) -0.01 1.23 1.61 0.72 - -
-BLCMV Beamformer [4] (2.827\3.304\3.495) -1.0 3.46 3.9 0.95 0.672 0.599
-BDE-BiTSE [15] (3.072\3.471\3.768) 24.01 3.66 4.02 0.99 0.142 0.485
-Bi-TSE-HRTF (proposed) (2.999\3.398\3.742) 20.18 3.14 3.78 0.98 0.1024 0.0042
-MLS-French Unprocessed (2.499\3.252\2.947) -0.01 1.23 1.63 0.71 - -
-BLCMV Beamformer (2.761\3.277\3.384) -1.03 3.37 3.82 0.94 0.694 0.608
-BDE-BiTSE (2.987\3.403\3.709) 23.67 3.67 4.04 0.99 0.3 0.503
-Bi-TSE-HRTF (proposed) (2.900\3.324\3.655) 19.86 3.14 3.79 0.98 0.1039 0.0041
 WSJ0 Unprocessed (2.845\3.376\3.411) 0 1.25 1.57 0.72 - -
@@
 Bi-TSE-HRTF-RI (3.110\3.504\3.764) 21.39 2.93 3.57 0.99 0.085 0.0044
-Bi-TSE-HRTF (proposed) (3.18\3.501\3.911) 21.41 3.47 3.88 0.99 0.0035 0.0756
+Bi-TSE-HRTF (proposed) (3.18\3.501\3.911) 21.41 3.47 3.88 0.99 0.0756 0.0035
 WSJ0 (Reverb) Unprocessed (2.348\3.069\2.783) -1.9 1.15 1.44 0.69 - -
@@
 Bi-TSE-HRTF (proposed) (2.86\3.348\3.481) 10.67 2.17 2.79 0.94 0.186 0.0082
-s
 […]
@@
 Bi-TSE-HRTF 0.0 -60 0
-4.4 Competing Methods
+IV-D Competing Methods
 To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear).
 Accordingly, we compared our method against the BLCMV beamformer [4] and a recent DNN-based model [15], both specifically designed for this scenario.
-The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the same training data as ours.
+The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the exact same training data as ours.
 We ensured that the strongest variant reported in the paper (BDE+CDF+IPD+SDF) was used for comparison.
-In addition, we trained two variants of our model.
-The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are split into their RI components and concatenated, resulting in four input channels instead of two.
+As an ablation study, we trained two variants of our model.
+The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are decomposed into their RI components and concatenated, yielding four input channels rather than two.
 The same decomposition is applied to the HRTF clue.
-The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue instead of the HRTF.
-In this case, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs, and passed through a fully connected layer to match the dimensionality of the bottleneck.
-For both variants, the loss function in (5) remains unchanged, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
-4.5 Results
-In Table 1, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions.
-In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE.
-Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
-Both DNN-based methods show strong cross-dataset generalization: though trained only on WSJ0-CSR1, they perform well on other corpora, including French, as shown in Table 1.
-We attribute this to the absence of speaker-specific priors and the reliance on a known HRTF database.
-Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI.
-The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation.
-By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation.
-Figure 3 shows the joint ITDILD p.d.f.s across methods.
-The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues.
+The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue rather than the HRTF.
+Here, the DOA is represented as a one-hot-encoded vector with the same resolution as the available HRTFs and passed through a fully connected layer to match the bottleneck’s dimensionality.
+For both variants, the loss function (5) remains unaltered, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
+IV-E Results
+[…]
+Figure 3 shows the joint ITDILD p.d.f.s for the extracted and target signals.
+The close alignment of the proposed method with the target distribution confirms effective preservation of binaural cues.
 In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations.
-The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f.
-becomes more concentrated, indicating dominance of the direct path in the output of the proposed method.
-Overall, it is evident that the proposed approach achieves this more effectively than competing techniques.
-Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig.
-3.
-Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec.
-4.3.
-The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
-Fig.
-4: A polar plot of SI-SDR versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue).
-Stars mark true DOAs; the dotted arc denotes the 0,dB threshold.
-Finally, to quantify spatial resolution, we use a “scanning” procedure: signals are extracted with HRTFs at azimuths in $90^{\circ},90^{\circ}]$, and the SI-SDR of each is computed against the target.
+The increased concentration of the joint distribution further indicates a dereverberation effect, with the direct path dominating the output.
+Figure 4: A polar plot of SI-SDR [dB] versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue). Stars mark true DOAs;
+Additionally, we evaluate two model variants: Bi-TSE-DOA and Bi-TSE-HRTF-RI.
+The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues—especially under reverberation—highlighting the central role of HRTFs in our formulation.
+The RI model achieves SI-SDR and binaural cue preservation comparable to those of the complex-valued model in anechoic conditions, but yields lower PESQ scores and degraded cue preservation under reverberation.
+To demonstrate spatial resolution, we employ a “scanning” procedure in which signals are extracted using HRTFs over azimuths in $90^{\circ},90^{\circ}]$, and the corresponding SI-SDR is computed with respect to the target.
 Plotting these values yields a beam-pattern-like diagram (Fig.
-4) with a clear peak at the true DOA.
-The main lobe width characterizes the model’s stability when the HRTF clue deviates mildly from the ground truth.
-5 Conclusions
-This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, yielding a personalized solution to the task at hand.
-Through cross-dataset evaluations, we demonstrate robust generalization across diverse corpora.
-The proposed approach not only produces high-quality speech but also faithfully preserves the binaural characteristics of the signal.
-Comparative analyses with existing binaural TSE methods reveal substantial improvements in spatial consistency, while preserving high speech quality.
-These findings highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction under a wide range of acoustic conditions.
+4) with a peak at the true DOA, where the main-lobe width reflects the model’s stability to mild HRTF deviations.
+V Conclusions
+This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, providing a personalized solution.
+By conditioning the extraction process on individualized HRTFs, the proposed approach preserves binaural structure while maintaining speech quality.
+Comparative analyses with existing binaural TSE methods demonstrate substantial improvements in spatial consistency, while preserving high speech quality.
+These results highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction across diverse acoustic conditions.